Ball A and B have a mass of 3 and 4 kg, respectively. If they have the velocities shown, d e 0.5. Determine their speeds just after direct central impact and the percentage of total (A-2 m/s (UB) 5 m/s energy loss for the two balls. A. VA-0.5 m/s, vB-4.0 m/s, AT-65% B. v,-4.0 m/s, vB-0.5 m/s, AT-65% C. VA-0.5 m/s, vB-4.0 m/s. ΔΤ-56% D. v,-4.0 m/s, vB 0.5 m/s. ΔΤ-56%
Answer #1

Let the final velocity be vA2 and vB2.

By momentum conservation,

mA*vA1 - mB*vB1 = mA*vA2 + mB*vB2

3*2 - 4*5 = 3vA2 + 4 vB2

Also, [vA1+ vB1] = (vB2 - vA2)/e

[5+2]*0.5 = vB2 - vA2

vA2 = -4 , vB2 = -0.5 m/s

speeds has to be in positive, vA = 4 m/s , vB = 0.5 m/s

total energy loss = 100*( 1 - [3*4^2 +4*0.5^2]/[3*2^2 +4*5^2]) = 56%

option d is correct.